OpenStudy (anonymous):

e^x-4e^-x=3 how would i solve this equation?

OpenStudy (campbell_st):

so the equation is \[e^x - \frac{4}{e^x} = 3\] multiply every term by \[e^x\] then you get \[e^{2x} - 4 = 3e^x\] or \[e^{2x} -3x^x - 4 = 0\] now you have an equation that can be reduced to a quadratic... by letting \[e^x = u\] so you have \[u^2 - 3u - 4 = 0\] solve for u, then make the reverse substitution and then solve for x. hope it helps

OpenStudy (anonymous):

@campbell_st thanks , so would the answer be u=4?

OpenStudy (campbell_st):

well you factor and have (u + 1)(u - 4) = 0 so u = 4 or -1 then the reverse substitution \[e^x = 4 ~~~and~~~e^x = -1\] so you now need to decide on the solution, and you have made a sound decision based on what you have posted. so now solve for x

OpenStudy (anonymous):

So then would you just substitute 4 in ? 4-4/4=3?

OpenStudy (anonymous):

\[e^{-1}\] has no solution, whereas if \(e^x=4\) then \(x=\ln(4)\)